The cost of hiring a catering service to serve food for a party is 150 pesos per head for 20 persons or less,130 pesos per head for 21 to 50 persons and
110 pesos per head for 51 to 100 persons.For 101 or more persons,the cost is at 100 pesos per head.Represent the total cost as a piecewise function of the number or attendances of the party.

Answer by Theo(13342)

cost is:

150 pesos per head for 20 persons or less.
130 pesos for 21 to 50 persons.
110 pesos per head for 51 to 100 persons.
100 pesos per head for 101 or more pesos.

let f(x) represent the cost function.

let x represent the number of persons.

your piece-wise function would look like:

f(x) = 150 * x for 0 <= x <= 20
f(x) = 130 * x for 21 <= x <= 50
f(x) = 110 * x for 51 <= x <= 100
f(x) = 100 * x for x >= 101

alternatively, it could also be written as:

f(x) = 0 for x = 0
f(x) = 150 * x for 0 < x <= 20
f(x) = 130 * x for 20 < x <= 50
f(x) = 110 * x for 50 < x <= 100
f(x) = 100 * x for 100 < x <= infinity

there are various ways to express it, but the main idea is that the function f(x) is defined differently for different intervals of x.

here's a reference:

https://www.mathsisfun.com/sets/functions-piecewise.html